How much work must be done on a system to decrease its volume from 13.0 L to 4.0 L by exerting a constant pressure of 5.0 atm?
Answer #1

Pressure, P-5.0 atm Change in volume, AV =(40-130)L =_ 9.0L work done, dw=-PxAV --(5.0 atm)× (-9.0 L) = 45.0 L . atm 101.3J し